# A Survey on Deep Learning for Theorem Proving

### v1

@@ 3.5 Other Tasks
In addition to the primary tasks outlined previously, we briefly list other prediction tasks related or helpful to theorem proving. A line of research (Urban & Jakubüv, 2020; Piotrowski & Urban, 2020b; Rabe et al., 2021; Johansson & Smallbone, 2023) explore automated conjecturing. Lee et al. (2020); Wu & Wu (2021) investigate proving theorems in the latent space. IsarStep (Li et al., 2021a) predicts the intermediate proposition given surrounding proofs. Skip-tree (Rabe et al., 2021) and PACT (Han et al., 2022) propose several self-supervised tasks by masking various proof terms for training language models. LIME (Wu et al., 2021c) constructs three reasoning tasks for deduction, abduction, and induction. Recently, Li et al. (2023) proposes to match the proofs with theorem statements from a large database and REFACTOR (Zhou et al., 2024b) predicts to extract theorems from proofs.
@@ 4 Datasets
This section classifies and summarizes datasets for theorem proving into 2 categories: i) datasets extracted from existing corpora or manually curated and ii) those using synthetic generation or augmentation methods. The overview of these datasets is shown in Figure 3.
@@ 4.1 Data Collection
We begin with the review of informal datasets. NL-PS (Ferreira & Freitas, 2020a) first builds a natural language premise selection dataset source from ProofWiki. Similarly, NaturalProofs (Welleck et al., 2021) further incorporates data from Stacks and textbooks, resulting in a dataset with roughly 25k examples. Adapted from it, NaturalProofs-Gen (Welleck et al., 2022) contains around 14.5k theorems for informal proof generation. Moreover, MATcH (Li et al., 2023) constructs over 180k statement-proof pairs for matching using the MREC corpus.
@@
For formal datasets, a line of efforts focuses on extracting and cleaning theorems and proofs from established formal libraries and verification projects. Notable datasets for Coq include Gamepad (Huang et al., 2019), CoqGym (Yang & Deng, 2019), and PRISM (Reichel et al., 2023), with CoqGym constructing a dataset from 123 projects encompassing 71k proofs. For Isabelle, datasets like IsarStep (Li et al., 2021a), PISA (Jiang et al., 2021), and Magnushammer (Mikuła et al., 2024) are built on the Archive of Formal Proofs and Isabelle Standard Library, where PISA extracts 183K theorems and 2.16M proof steps. LeanStep (Han et al., 2022), LeanDojo (Yang et al., 2023), and MLFMF (Bauer et al., 2023) leverage the mathlib library (mathlib Community, 2020) in Lean. In particular, LeanDojo extracts over 98k theorems and proofs with 130k premises from mathlib. Datasets for other proof assistants include HolStep (Kaliszyk et al., 2017) and HOList (Bansal et al., 2019) for HOL Light, MPTP2078 (Alama et al., 2014), Mizar40 (Kaliszyk & Urban, 2015b), and M2K (Kaliszyk et al., 2018) for Mizar, etc. Besides extracting data from existing projects, several works manually annotate or formalize the problems in natural language. Notably, MiniF2F (Zheng et al., 2022) manually formalizes 488 Olympiad-level problems across 4 proof systems and equally splits them into a validation set and a test set. FIMO (Liu et al., 2023) and ProofNet (Azerbayev et al., 2023) formalize the theorem statements of IMO and undergraduate-level problems in Lean. For other domains, TRIGO (Xiong et al., 2023) targets formalizing the trigonometric reduction problem, while UniGeo (Chen et al., 2022) and FormalGeo (Zhang et al., 2023c) annotate the proof steps for geometry proving problems.
@@
On the other hand, a large body of recent studies leverages large-scale online corpora with billions of tokens from informal and formal mathematical data to build the pre-training datasets that could aid in theorem proving. These datasets include WebMath (Han et al., 2022), Proof-Pile (Azerbayev et al., 2023), InternLM-Math (Ying et al., 2024), etc.
{forest}
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=black, rounded corners, align=left, minimum width=2.5em, edge+=black, line width=0.5pt, s sep=3pt, inner xsep=1.9pt, inner ysep=1.9pt, line width=0.5pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=0text width=2.2em,font=,, where level=1text width=4.4em,font=,, where level=2text width=3.8em,font=,, where level=3text width=38.6em,font=,, [ Datasets [Data Collection [Natural
Language[ E.g., NL-PS (Ferreira & Freitas, 2020a), NaturalProofs (Welleck et al., 2021), NaturalProofs-Gen (Welleck et al., 2022), MATcH (Li et al., 2023)]] [Coq [ E.g., GamePad (Huang et al., 2019), CoqGym (Yang & Deng, 2019), PRISM (Reichel et al., 2023)]] [Isabelle [ E.g., IsarStep (Li et al., 2021a), PISA (Jiang et al., 2021), MiniF2F (Zheng et al., 2022), Magnushammer (Mikuła et al., 2024)]] [Lean [ E.g., LeanStep (Han et al., 2022), MiniF2F (Zheng et al., 2022), FIMO (Liu et al., 2023), TRIGO (Xiong et al., 2023),
ProofNet (Azerbayev et al., 2023), LeanDojo (Yang et al., 2023), MLFMF (Bauer et al., 2023)]] [HOL Light [ E.g., HolStep (Kaliszyk et al., 2017), HOList (Bansal et al., 2019), MiniF2F (Zheng et al., 2022)]] [Mizar [ E.g., MPTP2078 (Alama et al., 2014), Mizar40 (Kaliszyk & Urban, 2015b), M2K (Kaliszyk et al., 2018)]] [Geometry [ E.g., UniGeo (Chen et al., 2022), FormalGeo (Zhang et al., 2023c)]] [Other
Languages [ E.g., Holophrasm (Whalen, 2016), TPTP (Sutcliffe, 2017), TacticToe (Gauthier et al., 2020), MLFMF (Bauer et al., 2023)]] [Pre-training
Corpus [ E.g., WebMath (Polu & Sutskever, 2020), Proof-Pile (Azerbayev et al., 2023), MathPile (Wang et al., 2023b), OpenWebMath (Paster et al., 2024),
Proof-Pile-v2 (Azerbayev et al., 2024), DeepSeekMath (Shao et al., 2024), InternLM-Math (Ying et al., 2024)]] ] [Data Generation [Rule-based
Generator [ E.g., INT (Wu et al., 2021b), MetaGen (Wang & Deng, 2020), LIME (Wu et al., 2021c), FwdP (Firoiu et al., 2021)
AutoTrig (Liu et al., 2022b), TRIGO-gen (Xiong et al., 2023), GeomVerse (Kazemi et al., 2023), AlphaGeometry (Trinh et al., 2024)]] [Iterative
Discovery [ E.g., LEGO-Prover (Xin et al., 2024), REFACTOR (Zhou et al., 2024b)]] [Other
Methods [ E.g., MMA (Jiang et al., 2023a), MUSTARD (Huang et al., 2024)]] ] ]
Figure 3: The taxonomy of datasets in theorem proving.
@@ 4.2 Data Generation
Beyond utilizing existing projects, researchers also study the generation of new proof data. A line of work (Wu et al., 2021b; c; Firoiu et al., 2021; Liu et al., 2022b; Xiong et al., 2023; Kazemi et al., 2023; Trinh et al., 2024) develops rule-based generators to generate both theorems and proofs by iteratively sampling inference rules and axioms from a pre-defined set to form a new theorem. Such a generation process enables manual control of difficulty. For example, INT (Wu et al., 2021b) synthesizes theorems for inequalities with different proof lengths while GeomVerse (Kazemi et al., 2023) targets geometry problems with 5 dimensions of difficulty design. MetaGen (Wang & Deng, 2020) further trains a neural generator to synthesize theorems similar to human-write ones.
Alternative approaches turn to iteratively augment the training dataset with fixed theorems but newly generated proofs. A line of work (Bansal et al., 2019; Polu & Sutskever, 2020; Polu et al., 2023) adopts the idea of expert iteration (Silver et al., 2018), which repeatedly applies the trained prover on existing theorems and adds the successful proof paths as new data points to train the prover further. Aygün et al. (2022) further proposes to adapt hindsight experience replay (Aygün et al., 2022) to FOL provers, which leverages previously unsuccessful proof trajectories by viewing their final states as the desired ones. Meanwhile, a line of reinforcement learning methods can also be viewed in this category.
Additionally, some works aim to generate intermediate helpful lemmas. REFACTOR (Zhou et al., 2024b) trains a GNN to extract lemmas from proofs, which can be used to streamline the proofs of other theorems. LEGO-Prover (Xin et al., 2024) prompts GPT-4 to generate sub-goal lemmas for the proof of a theorem and finalize it by proving or retrieving these lemmas. Newly proven lemmas are added to a growing library for future use. Works on conjecturing are also relevant to this field despite the generated conjectures may be incorrect.
Moreover, several recent approaches leverage auto(in)formalization for data generation. Using GPT-4, MMA (Jiang et al., 2023a) informalizes all theorem statements in Archive of Formal Proofs and mathlib, and MUSTARD (Huang et al., 2024) synthesizes problems from seed concepts, crafts informal proofs, and translates them into Lean to verify the correctness.
@@ 5 Evaluations
In this section, we focus on the evaluations of deep learning approaches in theorem proving, analyzing the key metrics and state-of-the-art performance for each task.
@@
Autoformalization. The assessment of autoformalization mainly relies on manually checking the equivalence between informal and formalized statements. Through such evaluation, recent studies (Wu et al., 2022; Azerbayev et al., 2023) show that state-of-the-art LLMs with few-shot prompting can correctly formalize 25% and 13% high-school and undergraduate-level problems, demonstrating the challenge in autoformalization. In addition, both studies reveal that LLMs exhibit considerably higher efficacy in informalization, achieving around 76% and 62% accuracies. Despite the modest success in autoformalizing statements, subsequent research (Jiang et al., 2023b; Zhao et al., 2023; Xin et al., 2024) reveals the value of autoformalizing informal intermediate goals to prove theorems in modularized pipelines.
@@
Premise selection. Retrieval metrics are widely used as the evaluation metric for premise selection. For example, recall at k (R@k) measures the ratio of correctly used premises in ground-truth proof within the top-k selections, and the mean reciprocal rank (MRR) computes the average reciprocal rank of the first correctly selected premise. Dense passage retrieval (DPR) (Karpukhin et al., 2020) with a Transformer encoder has significantly improved upon traditional methods, demonstrating a remarkable generalization ability to unseen data for premise selection. For example, ReProver (Yang et al., 2023) achieves 27.6% for R@10 and 0.24 for MRR on the LeanDojo benchmark for retrieving unseen premises in training, while BM25 (Robertson et al., 2009) achieves 15.5% for R@10 and 0.14 for MRR. Furthermore, a DPR-based retriever, Magnushammer (Mikuła et al., 2024), outperforms a symbolic method, Sledgehammer (Böhme & Nipkow, 2010), when used to select premises for the Thor prover (Jiang et al., 2022), improving the theorem proving success rate from 57% to 71% on the PISA dataset and from 28.3% to 36.9% on the MiniF2F-valid dataset.
@@
Theorem proving. The effectiveness of proofstep generation, proof search, and support from autoformalization and premise selection can be collectively evaluated by their success rate in proving theorems within a test set. Recent research (Zheng et al., 2023; Xin et al., 2024) shows impressive performance increases using state-of-the-art LLMs like GPT-4 in structured frameworks than tactic-based models with the best-first search. For instance, LEGO-Prover (Xin et al., 2024) achieves 57.0% accuracy and 50.0% accuracy on the valid and test set of MiniF2F, while the previous best model Thor (Jiang et al., 2022) with expert iteration training (Wu et al., 2022) achieves 37.3% and 35.2%. Moreover, deep learning-based proof searches could further improve the performance of ITP/ATP systems. HTPS (Lample et al., 2022) achieves an accumulative successful rate of 58.6% on MiniF2F-valid through online training and a 41.0% success rate with 64 search attempts on MiniF2F-test. Besides, the reinforcement learning-based ATP prover NIAGRA (Fokoue et al., 2023) outperforms both E (Schulz, 2002) and Vampire (Kovács & Voronkov, 2013) on the MPTP2078 dataset.
@@
Caveats. Tasks related to theorem proving can be tricky to evaluate. For autoformalization, manual evaluation is costly, whereas automated metrics are inaccurate. For example, the compilation rate evaluates only syntactic correctness, and the BLEU score (Papineni et al., 2002) struggles with formalizations semantically similar but not logically equivalent to the ground truth. For premise selection, metrics rely on the premises that are used in ground-truth proofs, so they may neglect other valid premises for alternative correct proofs different from the ground-truth ones, causing false negatives. Lastly, the evaluation of theorem proving is complicated by various experimental setups (§ 6.1).

### v2

@@ 3.5 Other Tasks
In addition to the primary tasks outlined previously, we briefly list several other prediction tasks that are related or helpful to theorem proving. One prominent line of research (Clark et al., 2020; Saha et al., 2020; Dalvi et al., 2021; Tafjord et al., 2021; Sanyal et al., 2022; Bostrom et al., 2022; Hong et al., 2022; Mishra et al., 2022; Yang et al., 2022; Tafjord et al., 2022; Morishita et al., 2023; Saparov & He, 2023) focuses on generating step-by-step rationale of a hypothesis from a set of known facts, which answers an open-domain question. These works primarily perform FOL rule deduction over natural language, in the form of an entailment tree, an analogy to the proof tree in formal theorem proving. Another area of interest (Urban & Jakubüv, 2020; Piotrowski & Urban, 2020b; Rabe et al., 2021; Johansson & Smallbone, 2023; Bengio & Malkin, 2024; Poesia et al., 2024) is automated conjecturing, which aims to discover new and interesting theorems beyond existing data. Additionally, research efforts such as Huang et al. (2019); Glorot et al. (2019); Polu et al. (2023) predict the proof length remained for a goal. Lee et al. (2020); Wu & Wu (2021) investigate proving theorems in the latent space. IsarStep (Li et al., 2021a) predicts the intermediate proposition given surrounding proofs. Skip-tree (Rabe et al., 2021) and PACT (Han et al., 2022) propose several self-supervised tasks by masking various proof terms for training language models. LIME (Wu et al., 2021c) creates three synthetic pre-training tasks inspired by three reasoning primitives of deduction, induction, and abduction, to improve the performance of formal theorem proving. Recently, Li et al. (2023) proposes to match the informal proofs with theorem statements from a large database, REFACTOR (Zhou et al., 2024b) and ATG (Lin et al., 2024c) focus on extracting or generating new useful formal theorems from proofs.
@@ 4 Datasets
This section classifies and summarizes datasets for theorem proving into 2 categories: i) datasets extracted from existing corpora or manually curated and ii) those using synthetic generation or augmentation methods. The taxonomy of these datasets is shown in Figure 3, and Table 1 in the Appendix provides more detailed information.
{forest}
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=black, rounded corners, align=left, minimum width=2.5em, edge+=black, line width=0.5pt, s sep=3pt, inner xsep=1.9pt, inner ysep=1.9pt, line width=0.5pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=0text width=2.2em,font=,, where level=1text width=4.4em,font=,, where level=2text width=3.8em,font=,, where level=3text width=39.4em,font=,, [ Datasets [Data Collection [Natural
Language[ E.g., NL-PS (Ferreira & Freitas, 2020a), NaturalProofs (Welleck et al., 2021), NaturalProofs-Gen (Welleck et al., 2022)]] [Coq [ E.g., GamePad (Huang et al., 2019), CoqGym (Yang & Deng, 2019), Proverbot9001 (Sanchez-Stern et al., 2020), PRISM (Reichel et al., 2023),
Graph2Tac (Rute et al., 2024)]] [Isabelle [ E.g., IsarStep (Li et al., 2021a), PISA (Jiang et al., 2021), miniF2F (Zheng et al., 2022), MAPL (Mikuła et al., 2024), Selene (Zhang et al., 2024),
FVELer (Lin et al., 2024b), PutnamBench (Tsoukalas et al., 2024)]] [Lean [ E.g., LeanStep (Han et al., 2022), miniF2F (Zheng et al., 2022), FIMO (Liu et al., 2023), TRIGO (Xiong et al., 2023), LeanDojo (Yang et al., 2023),
ProofNet (Azerbayev et al., 2023), miniCodeProps (Lohn & Welleck, 2024), PutnamBench (Tsoukalas et al., 2024), LEAN-GitHub (Wu et al., 2024)]] [HOL Light [ E.g., HolStep (Kaliszyk et al., 2017), HOList (Bansal et al., 2019), miniF2F (Zheng et al., 2022)]] [Mizar [ E.g., MPTP2078 (Alama et al., 2014), Mizar40 (Kaliszyk & Urban, 2015b), M2K (Kaliszyk et al., 2018)]] [Other
Languages [ E.g., TPTP (Sutcliffe, 2017), TacticToe (Gauthier et al., 2020), set.mm (Lample et al., 2022), MLFMF (Bauer et al., 2023), UniGeo (Chen et al., 2022),
IMO-AG-30 (Trinh et al., 2024)]] [Pre-training
Corpora [ E.g., WebMath (Polu & Sutskever, 2020), Proof-Pile (Azerbayev et al., 2023), MathPile (Wang et al., 2023b), OpenWebMath (Paster et al., 2024),
Proof-Pile-v2 (Azerbayev et al., 2024), DeepSeekMath (Shao et al., 2024), InternLM-Math (Ying et al., 2024b)]] ] [Data Generation [Rule-based
Generators [ E.g., INT (Wu et al., 2021b), MetaGen (Wang & Deng, 2020), LIME (Wu et al., 2021c), FwdP (Firoiu et al., 2021), PropL (An et al., 2024),
AlphaGeometry (Trinh et al., 2024), AIPS (Wei et al., 2024)]] [Iterative
Discovery [ E.g., LEGO-Prover (Xin et al., 2024b), REFACTOR (Zhou et al., 2024b), ATG (Lin et al., 2024c)]] [Auto(in)-
formalization [ E.g., MMA (Jiang et al., 2023a), FormL4 (Lu et al., 2024), MUSTARD (Huang et al., 2024), DeepSeek-Prover (Xin et al., 2024a),
Lean Workbook (Ying et al., 2024a)]] ] ]
Figure 3: The taxonomy of datasets in theorem proving.
@@ 4.1 Data Collection
We begin with the review of informal datasets. NL-PS (Ferreira & Freitas, 2020a) first builds a natural language premise selection dataset source from ProofWiki. Similarly, NaturalProofs (Welleck et al., 2021) further incorporates theorems from Stacks, and textbooks for premise selection. Adapted from NaturalProofs, NaturalProofs-Gen (Welleck et al., 2022) utilizes a subset of theorems and proofs for informal proof generation.
@@
For formal datasets, a line of efforts has been made to extract and clean theorems and proofs from established formal libraries and verification projects. Notable datasets for Coq include GamePad (Huang et al., 2019), CoqGym (Yang & Deng, 2019), Proverbot9001 (Sanchez-Stern et al., 2020), PRISM (Reichel et al., 2023), and Graph2Tac (Rute et al., 2024), which are constructed based on mathematical or software verification projects. For Isabelle, datasets like IsarStep (Li et al., 2021a), PISA (Jiang et al., 2021), and MAPL (Mikuła et al., 2024) are built on the Archive of Formal Proofs and Isabelle Standard Library, while Selene (Zhang et al., 2024) and FVELer (Lin et al., 2024b) are constructed based on the seL4 project (Klein et al., 2009). LeanStep (Han et al., 2022), LeanDojo (Yang et al., 2023), MLFMF (Bauer et al., 2023), and LEAN-GitHub (Wu et al., 2024) utilize Lean’s open-source libraries (e.g., the mathlib library (mathlib Community, 2020)). Datasets for other proof assistants include HolStep (Kaliszyk et al., 2017) and HOList (Bansal et al., 2019) for HOL Light, MPTP2078 (Alama et al., 2014), Mizar40 (Kaliszyk & Urban, 2015b), and M2K (Kaliszyk et al., 2018) for Mizar, etc. Besides extracting data from existing projects, several works manually formalize or annotate the problems into formal languages: miniF2F (Zheng et al., 2022), FIMO (Liu et al., 2023), ProofNet (Azerbayev et al., 2023), and PutnamBench (Tsoukalas et al., 2024) manually formalize high school and college-level problems in mathematical competitions or textbooks in Lean or Isabelle, and miniCodeProps (Lohn & Welleck, 2024) translates Haskell programs into Lean. For other domains, TRIGO (Xiong et al., 2023) formalizes the trigonometric reduction problem in Lean, while UniGeo (Chen et al., 2022) and IMO-AG-30 (Trinh et al., 2024) annotate proof steps for geometry proving problems in their designed formal languages.
@@
On the other hand, a large body of recent studies leverages large-scale online corpora with billions of tokens from informal and formal mathematical data to build the pre-training datasets that could aid in theorem proving. These datasets include WebMath (Polu & Sutskever, 2020), Proof-Pile (Azerbayev et al., 2023), DeepSeekMath (Shao et al., 2024), etc.
@@ 4.2 Data Generation
Beyond utilizing existing projects, researchers also study the generation of new theorems and proofs. A line of work (Wu et al., 2021b; c; Firoiu et al., 2021; Trinh et al., 2024; An et al., 2024; Wei et al., 2024) develops rule-based generators to produce new data by iteratively sampling inference rules from a pre-defined set. This process enables manual control over both the quantity and difficulty of the generated theorems. For example, INT (Wu et al., 2021b) produces 1.5 million inequality training problems with different proof lengths and axiom distributions than the testing ones, while AlphaGeometry (Trinh et al., 2024) synthesizes over 100 million training data with proof lengths ranging from 1 to 247. MetaGen (Wang & Deng, 2020) also trains a neural generator to produce theorems similar to human-write ones.
Alternative approaches turn to iteratively augment the training dataset with fixed theorems but newly generated proofs. A line of work (Bansal et al., 2019; Polu & Sutskever, 2020; Polu et al., 2023) adopts the idea of expert iteration (Silver et al., 2018), which repeatedly applies the trained prover on existing theorems and adds the successful proof paths as new data points to train the prover further. Aygün et al. (2022) also proposes to adapt hindsight experience replay (Andrychowicz et al., 2017) to FOL provers, which leverages previously unsuccessful proof trajectories by viewing their final states as the desired ones. Meanwhile, a line of RL-based methods can also be viewed in this category.
Additionally, some studies aim to generate intermediate helpful lemmas. REFACTOR (Zhou et al., 2024b) trains a GNN to extract lemmas from proofs, which can be used to streamline the proofs of other theorems. Similarly, ATG (Lin et al., 2024c) generates new theorems based on MCTS and self-play learning to shorten proofs. LEGO-Prover (Xin et al., 2024b) prompts GPT-4 to generate sub-goal lemmas for the proof of a theorem and finalize it by proving or retrieving these lemmas. Newly proven lemmas are added to a library for future use. Works on conjecturing are also relevant to this field, although the generated conjectures could be more challenging than existing theorems and may be incorrect or difficult to prove.
Moreover, several recent approaches leverage auto(in)formalization for data generation. Using GPT-4, MMA (Jiang et al., 2023a) informalizes all theorem statements in Archive of Formal Proofs and mathlib, while FormL4 (Lu et al., 2024) generates informal descriptions of both theorems and proofs extracted from mathlib. MUSTARD (Huang et al., 2024) synthesizes problems from a few sampled seed concepts, crafts informal proofs, and translates them into Lean to verify their correctness. DeepSeek-Prover (Xin et al., 2024a) autoformalizes mathematical competition problems and filters high-quality statements through model scoring and hypothesis rejection methods. Similarly, Lean Workbook (Ying et al., 2024a) proposes an active learning pipeline that autoformalizes mathematical questions and filters them through Lean’s compilation, LLMs’ evaluation, and human diagnostics at each round.
@@ 5 Evaluations
In this section, we focus on the evaluations of deep learning approaches in theorem proving, analyzing the key metrics and state-of-the-art performance for each task. Detailed accuracies of state-of-the-art methods on several key datasets are provided in Table 2 in the Appendix.
@@
Autoformalization. The assessment of autoformalization mainly relies on manually checking the equivalence between informal and formalized statements. Recent studies (Wu et al., 2022; Azerbayev et al., 2023) indicate that state-of-the-art LLMs with few-shot prompting can only correctly formalize 25% and 13% high-school and undergraduate-level problems, highlighting the challenge in autoformalization. In addition, both studies reveal that LLMs exhibit considerably higher efficacy in informalization, achieving accuracies of around 76% and 62%, respectively. Despite the modest success in autoformalizing statements, subsequent research (Jiang et al., 2023b; Zhao et al., 2024; Xin et al., 2024b) reveals the benefit of autoformalizing intermediate sub-goals to generate formal proofs in modularized pipelines.
@@
Premise Selection. Retrieval metrics are widely used as the evaluation metric for premise selection. For example, recall at $k$ (R@$k$) measures the ratio of correctly used premises in ground-truth proof within the top-$k$ selections, and the mean reciprocal rank (MRR) computes the average reciprocal rank of the first correctly selected premise. Among existing methods, DPR with a Transformer encoder has significantly improved upon traditional methods, demonstrating a remarkable generalization ability to unseen data for premise selection. For example, ReProver (Yang et al., 2023) achieves 27.6% for R@10 and 0.24 for MRR on the LeanDojo benchmark for retrieving unseen premises in training, while the baseline method BM25 (Robertson et al., 2009) achieves 15.5% for R@10 and 0.14 for MRR. Furthermore, a DPR-based retriever, Magnushammer (Mikuła et al., 2024), outperforms the classic hammer system, Sledgehammer (Böhme & Nipkow, 2010), when used to select premises for the Thor prover (Jiang et al., 2022), improving the theorem proving success rate from 57% to 71% on the PISA dataset and from 28.3% to 36.9% on the miniF2F-valid dataset.
@@
Theorem Proving. The effectiveness of proofstep generation, proof search, and support from autoformalization and premise selection can be collectively evaluated by their success rate in proving theorems within a test set. Recent research (Zheng et al., 2024; Xin et al., 2024b) shows impressive performance increases using state-of-the-art LLMs like GPT-4 in structured frameworks than fine-tuned tactic-based language models with search heuristics. For instance, LEGO-Prover (Xin et al., 2024b) achieves 57.0% accuracy and 50.0% accuracy on the valid and test set of miniF2F, while the previous best prover Thor (Jiang et al., 2022) with expert iteration training (Wu et al., 2022) achieves 37.3% and 35.2%. Moreover, advanced learning-based proof searches could further improve the performance of ITP/ATP systems. HTPS (Lample et al., 2022) achieves an accumulative successful rate of 58.6% on miniF2F-valid through online training and a 41.0% success rate with 64 search attempts on miniF2F-test. Besides, the RL-based ATP system NIAGRA (Fokoue et al., 2023) outperforms both E (Schulz, 2002) and Vampire (Kovács & Voronkov, 2013) on the MPTP2078 dataset.
@@
Caveats. Tasks related to theorem proving can be tricky to evaluate. For autoformalization, manual evaluation is costly, whereas most automated metrics are inaccurate. For example, the compilation rate evaluates only syntactic correctness, and the BLEU score (Papineni et al., 2002) struggles with formalizations semantically similar but not logically equivalent to the ground truth. Although LeanEuclid (Murphy et al., 2024) provides automatic semantic evaluation between the autoformalized theorem statement and the ground truth, its domain is limited to Euclidean geometry, making it difficult to generalize to other areas. For premise selection, metrics rely on the premises that are used in ground-truth proofs, so they may neglect other valid premises for alternative correct proofs different from the ground-truth ones, causing false negatives. Additionally, the evaluation of theorem proving is further complicated by the variety of experimental setups, as detailed in §6.1.